# Information Elicitation in Agency Games

### v2

@@ 3.2.2 Concealment and revelation conditions under a decreasing ratio assumption
Assumption 2 is analogous to the “increasing ratio condition” assumption from Aguirre et al. [2010], which instead has the derivative of total welfare in the numerator. Our analysis naturally extends this to focus on agent utility. Assumption 2 holds in almost the same set of conditions as the assumption on total welfare from Aguirre et al. [2010], and we discuss the subtleties of the differences between these assumptions in Appendix A.3.
@@
Under Assumptions 1 and 2, $V_{\text{const}}(r)$ is strictly quasi-convex for $r\in[0,p_{0}^{*}-p_{1}^{*}]$. That is, if there exists $\hat{r}\in[0,p_{0}^{*}-p_{1}^{*}]$ such that $V_{\text{const}}(\hat{r})=0$, then $V_{\text{const}}^{\prime\prime}(\hat{r})>0$.
@@
Under Assumptions 1 and 2, $V_{\text{con}}(p^{*})>V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ if
@@
Under Assumptions 1 and 2, $V_{\text{con}}(p^{*})<V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ if
@@
An important property analyzed by Milgrom [1981] is the monotone likelihood ratio property (MLRP), which here would say that $\frac{f_{0}(c)}{f_{1}(c)}$ is increasing for all $c$. The MLRP would imply that both sides of the inequality in equation (6) are greater than 1. However, this does not necessarily imply an order between these ratios, and there exist distributions that satisfy the MLRP that yield either of the inequality directions above. We give a specific example of this using the Weibull distribution in Section 3.3 below.
3.3 Example: Exponential and Weibull Distributions
@@
3.4 Principal’s Revelation Preferences
@@
3.5 Total Welfare Consequences
We now consider whether revealing $X$ increases total welfare, or the sum of utilities of the principal and agent. When $X$ is concealed, total welfare is given by
$W_{\text{con}}(p)\coloneqq V_{\text{con}}(p)+\Pi_{\text{con}}(p),$
and when $X$ is revealed, total welfare is given by
$W_{\text{rev}}(p_{0},p_{1})\coloneqq V_{\text{rev}}(p_{0},p_{1})+\Pi_{\text{rev}}(p_{0},p_{1}).$
The question of whether price discrimination increases total welfare has been well studied [Varian, 1985]. Our comparison of the concealed vs. revealed contracts has a direct isomorphism with third degree monopoly price discrimination. Specifically, our concealed setting corresponds to monopoly pricing without price discrimination, with the seller acting as principal and buyer acting as agent. Our revealed setting corresponds to a monopoly seller enacting third degree price discrimination over markets segmented by $X$.
Thus, with minor adjustments, we can apply results from the price discrimination literature that characterize the effects of third degree monopoly price discrimination on total welfare. Mirroring Varian [1985]’s seminal work, Lemma 4 shows that total welfare increases only if the quantity of tasks completed also increases in the revealed setting compared to the concealed setting.
Lemma 4.
Total welfare increases under revelation ($W_{\text{rev}}(p_{0}^{*},p_{1}^{*})-W_{\text{con}}(p^{*})$) only if task completion quantity increases under revelation,
$F(p^{*})<(1-\theta)F_{0}(p_{0}^{*})+\theta F_{1}(p_{1}^{*}).$
3.5.1 Example when total welfare decreases
@@
4 Information Revelation With Garbling
So far, we have compared the settings when the environmental variable $X$ is either concealed or and revealed, focusing on the agent’s incentive to induce each setting at time $t=1$. We now generalize the agent’s action space to instead be able to reveal a garbled version of the variable $X$. From a modeling standpoint, the agent’s garbling action space interpolates between the concealed and revealed settings. In this work, we consider a randomized response garbling mechanism which has been applied in many settings, from statistical informativeness [Blackwell, 1951] to survey experiment design [Warner, 1965] to differential privacy [Dwork et al., 2006, Kasiviswanathan et al., 2011].
As an overview of results in this section, we show that there exist conditions under which the agent would prefer to garble over both concealment and revelation. Thus, having the option to garble can benefit the agent, and even induce the agent to reveal more information when they would otherwise fully opt to conceal $X$. We also analyze the effects of garbling on the principal’s utility and total welfare.
4.1 Model Setup: Information Transfer with Garbling
Suppose the agent has the option to present the principal with a different variable $Y$ with random noise on top of $X$, defined as
$Y=\begin{cases}X&\text{w.p. }\varepsilon\\ \xi&\text{w.p. }1-\varepsilon\end{cases}=\begin{cases}X&\text{w.p. }\frac{1+\varepsilon}{2}\\ \neg X&\text{w.p. }\frac{1-\varepsilon}{2},\end{cases}$ (10)
where $\xi\sim\text{Bernoulli}(\gamma)$ is independent of $X$ and $C$. This work considers this randomized response mechanism for binary $X$, though further extensions with different noise models can be made for continuous $X$.
The game with garbling proceeds as before, but the agent selects $\varepsilon\in[0,1]$ at time $t=1$. The full timing is outlined in Figure 3. The agency game with garbled information transfer is a generalization of the previous game: in the previous timing in Figure 1, the agent’s choice at $t=1$ would be equivalent to selecting $\varepsilon$ from a more restricted set $\{0,1\}$.
$t=0$ P and A share prior distribution over $C$. A knows
prior joint
distribution
of $C,X$.
$t=1$ A selects $\varepsilon$,
and reveals the
observability of $Y$ (10), including joint distribution of $C,Y$.
$t=2$ P offers a contract, which depends on $Y$. $t=3$ $X$ and $Y$ are
realized.
A learns their cost type $C$.
It is assumed that the principal, on knowing $Y$, cannot back out the previous value of $X$. The principal treats the variable $Y$ as an environmental variable with prior joint distribution $Y,C$, and proceeds to design the optimal contract conditioning on $Y$. The principal does not have access to $\varepsilon$ when designing the contract—only the variable $Y$ and its potential values.
The game proceeds as delineated for the revealed information contract in Section 3.1.2, but using the variable $Y\in\mathcal{Y}$ instead of $X$. Specifically, the principal designs a contract with transfer function $\rho:\mathcal{Y}\to\mathbb{R}_{+}$, and upon execution of the contract, the agent receives transfer $\rho(y)$ if they exert effort and $Y=y$. For binary $X$ and $Y$ following equation (10), the principal chooses $\rho(0)=p_{0}$ and $\rho(1)=p_{1}$.
Let the conditional distribution of the cost $C$ given $Y$ be distributed with CDF ${\mathbb{P}}(C\leq c|Y=y)=G_{y}(c)$. We focus our analysis on a simple randomized response noise model that does not change the marginal distribution over $X$, setting $\gamma=\theta=\frac{1}{2}$. In this case, we have
$G_{0}(c)=\frac{1+\varepsilon}{2}F_{0}(c)+\frac{1-\varepsilon}{2}F_{1}(c),\quad\text{and}\quad G_{1}(c)=\frac{1+\varepsilon}{2}F_{1}(c)+\frac{1-\varepsilon}{2}F_{0}(c).$
Similar results can be derived for general $\gamma,\theta$, and there is further room to explore the interaction between these parameters in future work.
As in Section 3.1, the utility of the principal under revelation of a given garbled variable $Y$ defined by equation (10) is given by
$\Pi_{\text{garb}}(p_{0},p_{1})=(1-\theta)E[(b-p_{0})\operatorname{\mathbbm{1}}(C<p_{0})|Y=0]+\theta E[(b-p_{1})\operatorname{\mathbbm{1}}(C<p_{1})|Y=1],$
and the utility of the agent is given by
$V_{\text{garb}}(p_{0},p_{1})=(1-\theta)E[(p_{0}-C)\operatorname{\mathbbm{1}}(C<p_{0})|Y=0]+\theta E[(p_{1}-C)\operatorname{\mathbbm{1}}(C<p_{1})|Y=1].$
The principal moves first and chooses $p_{0}(\varepsilon),p_{1}(\varepsilon)\in\operatorname*{arg\,max}_{p_{0},p_{1}}\Pi_{\text{garb}}(p_{0},p_{1})$. The agent’s resulting utility is $V_{\text{garb}}(p_{0}(\varepsilon),p_{1}(\varepsilon))$.
To simplify notation, we make the following abbreviations in the rest of this section: $\Pi_{\text{garb}}(\varepsilon)=\Pi_{\text{garb}}(p_{0}(\varepsilon),p_{1}(\varepsilon))$, $V_{\text{garb}}(\varepsilon)=V_{\text{garb}}(p_{0}(\varepsilon),p_{1}(\varepsilon))$.
4.2 Garbling and Prices
Lemma 5 (Monotonic price changes).
Suppose $\gamma=\theta=\frac{1}{2}$. Suppose the principal’s utility is strictly concave as a function of price (Assumption 1). Suppose the MLRP holds (Assumption 3). Then $p_{0}^{\prime}(\varepsilon)>0$ and $p_{1}^{\prime}(\varepsilon)<0$ for all $\varepsilon\in[0,1]$.
4.3 Agent’s Garbling Incentives
For the agent, there is no clear ordering between $V_{\text{garb}}(p_{0}(\varepsilon),p_{1}(\varepsilon))$, $V_{\text{con}}(p^{*})$, and $V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$. Garbling can be better for the agent than concealment, even when hiding is preferred over revelation, $V_{\text{con}}(p^{*})>V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$. Garbling can also be better than revelation, even when revelation is preferred over hiding, $V_{\text{con}}(p^{*})<V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$. We illustrate both of these cases using exponential distributions in Section 4.4, as previously done for the full concealment and revelation comparison in Section 3.3.
To give a more general theoretical characterization of the agent’s garbling incentives, we give sufficient conditions for the agent to prefer a nonzero amount of garbling over full revelation. This can be seen as a softer version of the previous analysis of when the agent prefers full concealment over full revelation.
4.3.1 Garbling condition under one zero-cost type
As in Section 3.2.1, we first analyze the restricted case where one of the agent types is anchored at zero-cost: suppose $C|X=1$ takes value 00 with probability one. Proposition 4 gives a sufficient condition for the agent to prefer a non-zero amount of garbling over full revelation.
Proposition 4 (Sufficient garbling condition with zero-cost type).
$\frac{b-p_{0}^{*}}{2-\sigma_{0}(p_{0}^{*})}<g_{0}(p_{0}^{*}),$ (11)
[…]
Notably, the inequality in equation (11) captures distributions that are not captured by Proposition 1. Thus, comparing Proposition 4 to Proposition 1 shows that the agent might want to garble, even if they may not always want to hide. In fact, the condition in equation (11) is quite general, and we show in the example in Section 4.4 below that equation (11) applies to any log-concave Weibull distribution.
4.3.2 General garbling condition
Definition 1 (Agent utility dominance).
Proposition 5 (Sufficient garbling condition).
4.4 Example: Exponential and Weibull Distributions
The significance of this example is that if one agent type is anchored at 00, and the non-zero-cost environment induces an exponential distribution, the agent will always have an incentive to garble, regardless of the mean of the non-zero-cost distribution. Consider this in comparison to the exponential example from Section 3.3, where the condition in Proposition 1 showed that agent prefers to fully conceal $X$ when $\lambda_{0}$ is small enough.
Similarly to Figure 2, we can also consider simulations beyond the zero-cost anchored setting by considering all combinations of $\lambda_{0},\lambda_{1}$ when $C|X=0\sim\text{Exp}(\frac{1}{\lambda_{0}})$ and $C|X=1\sim\text{Exp}(\frac{1}{\lambda_{1}})$, with $F_{x}$ given by equation (8). Figure 4 illustrates the combinations of $\lambda_{0},\lambda_{1}$ for which the agent prefers some amount of garbling over full revelation. Figure 4 shows that $V_{\text{garb}}^{\prime}(1)<0$ as long one of the conditional means $\lambda_{0}$ or $\lambda_{1}$ is small enough. That is, as long as one of the revealed settings has low enough cost, the agent always prefers to garble, regardless of how high the cost of the other setting goes. This contrasts the revelation example in Section 3.3, where even if $\lambda_{1}$ is close to 00, high enough $\lambda_{0}$ leads to the agent being willing to reveal.
{tblr}
p0.1cmc &$V_{\text{garb}}^{\prime}(1)$
[…]
$\lambda_{0}$
4.5 Principal’s Garbling Preferences
Similarly to Section 3.4, the principal always prefers for more information to be revealed. In fact, the garbling parameter $\varepsilon$ directly interpolates between the concealed and revealed utilities for the principal.
Lemma 6.
Lemma 7 (Strict principal improvement).
4.6 Garbling and Total Welfare
We now discuss the relationship between the agent’s the choice of garbling amount $\varepsilon$ and total welfare, $W_{\text{garb}}(p_{0},p_{1})\coloneqq\Pi_{\text{garb}}(p_{0},p_{1})+V_{\text{garb}}(p_{0},p_{1})$. For notational convenience, let $W_{\text{garb}}(\varepsilon)\coloneqq W_{\text{garb}}(p_{0}(\varepsilon),p_{1}(\varepsilon))$.
First, we show that relative to the fully concealed setting, the marginal effect of revealing any information on total welfare is initially positive relative to full concealment.
Lemma 8 (More information initially increases total welfare).
$W_{\text{garb}}^{\prime}(0)\geq 0$. The inequality is strict if $\Pi_{\text{garb}}^{\prime}(0)>0$, which is true under strict concavity of $\Pi_{0}(p),\Pi_{1}(p)$ (Assumption 1) and the MLRP (Assumption 3).
Second, we also know that the optimal amount of garbling chosen by the agent also increases total welfare over the fully concealed setting.
Lemma 9 (Optimal garbling increases welfare over concealment).
Intuitively, Lemma 9 follows from the fact that the principal is never hurt by additional information. Given that the agent benefits from their optimal garbling choice, total welfare must increase.
The increase in total welfare is important from a mechanistic standpoint: consider a principal deciding whether or not to allow for agent garbling in an information transfer mechanism. There exist cases when the agent would reveal some amount of garbled information in the game, but would choose to conceal if only allowed to choose between concealment and revelation (see Figure 5 for an example). Thus, the garbling mechanism leads to higher total welfare in those situations.
While the agent’s chosen amount of garbling $\varepsilon^{*}$ improves total welfare over concealment, the question remains of where $\varepsilon^{*}$ falls relative to the optimal amount of garbling that maximizes total welfare. Lemma 10 shows that the optimal amount of garbling that maximizes total welfare must necessarily reveal at least as much information as the optimal amount of garbling chosen by the agent. Under Assumptions 1 and 3, if $\varepsilon^{*}<1$, then the $\varepsilon$ that maximizes total welfare is strictly higher than the $\varepsilon$ that maximizes agent utility.
4.7 Garbling vs. Restricted Price Discrimination
Garbling is not the only way to interpolate between the full concealment and full revelation settings. In the price discrimination literature, there is an established model that interpolates between full price discrimination and no price discrimination by restricting that price difference between market segments can be no greater than some parameter $r$. Wright [1993] models this restriction as arising from a “cost of transport” or arbitrage between two markets. Aguirre et al. [2010] apply this interpolation by analyzing the marginal effect of $r$ on total welfare. We refer to this interpolation using $r$ as a restricted price discrimination model. We also leveraged this technique to analyze the agent’s utility in Section 3.2.
Principal utility, $\Pi$ Refer to caption
Agent utility, $V$
5 Conclusions and Future Work
We have presented a model in which an agent decides whether or not to reveal the observability of an environmental variable to a principal. We first considered a simple action space where the agent must choose between concealing or revealing a given variable $X$, and we later relaxed this to a continuous action space where the agent can choose to reveal a garbled version of the given environmental variable $X$. In both cases, we gave sufficient conditions on the conditional cost distributions for the agent to prefer revelation over concealment, concealment over revelation, and garbling over revelation. Our agency game has a price discrimination analog, which makes it possible to leverage existing results from price discrimination to analyze total welfare; however, our model also offers a qualitatively different perspective and interpolation method via garbling.
More broadly, this work was motivated by analyzing a mechanism by which one might discover unknown unknowns. Even as data and computational methods become increasingly sophisticated and widely available, this problem of discovery of which metrics or variables to analyze continues to permeate the natural sciences, social sciences, and engineering. In this work, we’ve explored one avenue for discovery of metrics in a setting of information asymmetry where relevant environmental variables are unknown to a principal, but known to an evaluated agent. There are other possibilities for formulating the question of who holds relevant information, and when they would be willing to share it. For example, one could model incentives for third-party individuals to offer new metrics, or perhaps bi-directional information transfer where a principal and agent both hold distinct information. The key element of our work that may be worth retaining in alternative information design frameworks is the property that the variable itself may be unknown to the information receiver.
Going beyond information design, our broader motivation is to apply these insights to help design mechanisms that would benefit all players. For example, if sharing information would increase total welfare, but decrease an agent’s private utility, then perhaps a wealth transfer mechanism exists to adequately compensate the agent for revealing the information. Or, perhaps there exist auction or tournament frameworks that may further encourage information sharing. Finally, a particularly interesting avenue for future work would be to consider the possibility of cooperation and competition between multiple agents in this agency game with information transfer.
Acknowledgements
Appendix A Proofs from Section 3
Here we give full proofs from Section 3.
A.1 Proof of Proposition 1
@@ A.2 Proofs for Propositions 2 and 3
We first prove Lemma 1, which follows from Assumption 1 and 2.
@@
Under Assumptions 1 and 2, $V_{\text{const}}(r)$ is strictly quasi-convex for $r\in[0,p_{0}^{*}-p_{1}^{*}]$. That is, if there exists $\hat{r}\in[0,p_{0}^{*}-p_{1}^{*}]$ such that $V_{\text{const}}(\hat{r})=0$, then $V_{\text{const}}^{\prime\prime}(\hat{r})>0$.
@@
$\begin{split}V_{\text{const}}^{\prime}(r)&=(1-\theta)V_{0}(p_{0}(r))p_{0}^{\prime}(r)+\theta V_{1}(p_{1}(r))p_{1}^{\prime}(r)\\ &=\left(\frac{-\Pi_{1}^{\prime\prime}(p_{1}(r)\Pi_{0}^{\prime\prime}(p_{0}(r)))}{\Pi_{0}^{\prime\prime}(p_{0}(r))+\Pi_{1}^{\prime\prime}(p_{1}(r))}\right)\left(\frac{\theta V_{1}^{\prime}(p_{1}(r))}{\Pi_{1}^{\prime\prime}(p_{1}(r))}-\frac{(1-\theta)V_{0}^{\prime}(p_{0}(r))}{\Pi_{0}^{\prime\prime}(p_{0}(r))}\right)\\ &=\left(\frac{-\Pi_{1}^{\prime\prime}(p_{1}(r)\Pi_{0}^{\prime\prime}(p_{0}(r)))}{\Pi_{0}^{\prime\prime}(p_{0}(r))+\Pi_{1}^{\prime\prime}(p_{1}(r))}\right)\left(\theta w_{1}(p_{1}(r))-(1-\theta)w_{0}(p_{0}(r))\right)\\ \end{split}$ (16)
@@
Under Assumptions 1 and 2, $V_{\text{con}}(p^{*})>V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ if
@@
$\Pi^{\prime\prime}_{x}(p)=-2f_{x}(p)+f_{x}^{\prime}(p)(b-p)$
@@
Under Assumptions 1 and 2, $V_{\text{con}}(p^{*})<V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ if
@@
$\frac{\theta F_{1}(p^{*})/f_{1}(p^{*})}{2+L(p^{*})\alpha_{1}(p^{*})}<\frac{(1-\theta)F_{0}(p^{*})/f_{0}(p^{*})}{2+L(p^{*})\alpha_{0}(p^{*})}$
A.3 Comparison of decreasing ratio condition to increasing ratio condition
The results in Section 3.2.2 depend on the decreasing ratio condition (DRC) given in Assumption 2. This is analogous to the “increasing ratio condition (IRC)” from Aguirre et al. [2010], which says that the ratio $\frac{W_{x}^{\prime}(p)}{\Pi_{x}^{\prime\prime}(p)}$ is increasing in $p$. We now discuss in more detail the relationship between the DRC and the IRC, including sufficient conditions under which the DRC holds. In introducing the IRC, Aguirre et al. [2010] describe a “very large set of demand functions” for which the IRC holds. Aguirre et al. [2010] give sufficient conditions for the IRC to hold in Appendix B from their paper, which includes linear functions and exponential and constant elasticity functions.

### v3

@@ D.2.2 Concealment and revelation conditions under a decreasing ratio assumption
Assumption 5 is analogous to the “increasing ratio condition” assumption from Aguirre et al. [2010], which instead has the derivative of total welfare in the numerator. Our analysis naturally extends this to focus on agent utility. Assumption 5 holds in almost the same set of conditions as the assumption on total welfare from Aguirre et al. [2010], and we discuss the subtleties of the differences between these assumptions in Appendix F.3.
@@
Under Assumptions 4 and 5, $V_{\text{const}}(r)$ is strictly quasi-convex for $r\in[0,p_{0}^{*}-p_{1}^{*}]$. That is, if there exists $\hat{r}\in[0,p_{0}^{*}-p_{1}^{*}]$ such that $V_{\text{const}}(\hat{r})=0$, then $V_{\text{const}}^{\prime\prime}(\hat{r})>0$.
@@
Under Assumptions 4 and 5, $V_{\text{con}}(p^{*})>V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ if
@@
Under Assumptions 4 and 5, $V_{\text{con}}(p^{*})<V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ if
@@
An important property analyzed by Milgrom [1981] is the monotone likelihood ratio property (MLRP), which here would say that $\frac{f_{0}(c)}{f_{1}(c)}$ is increasing for all $c$. The MLRP would imply that both sides of the inequality in equation (10) are greater than 1. However, this does not necessarily imply an order between these ratios, and there exist distributions that satisfy the MLRP that yield either of the inequality directions above. We give a specific example of this using the Weibull distribution in Section D.3 below.
D.3 Example: Exponential and Weibull Distributions
@@
D.4 Principal’s Revelation Preferences
@@
Appendix E Example Where Total Welfare Decreases under Revelation
@@
Appendix F Proofs from Section D
Here we give full proofs from Section D.
F.1 Proof of Proposition 1
@@ F.2 Proofs for Propositions 2 and 3
We first prove Lemma 7, which follows from Assumption 4 and 5.
@@
Under Assumptions 4 and 5, $V_{\text{const}}(r)$ is strictly quasi-convex for $r\in[0,p_{0}^{*}-p_{1}^{*}]$. That is, if there exists $\hat{r}\in[0,p_{0}^{*}-p_{1}^{*}]$ such that $V_{\text{const}}(\hat{r})=0$, then $V_{\text{const}}^{\prime\prime}(\hat{r})>0$.
@@
$\begin{split}V_{\text{const}}^{\prime}(r)&=(1-\theta)V_{0}(p_{0}(r))p_{0}^{\prime}(r)+\theta V_{1}(p_{1}(r))p_{1}^{\prime}(r)\\ &=\left(\frac{-\Pi_{1}^{\prime\prime}(p_{1}(r)\Pi_{0}^{\prime\prime}(p_{0}(r)))}{\Pi_{0}^{\prime\prime}(p_{0}(r))+\Pi_{1}^{\prime\prime}(p_{1}(r))}\right)\left(\frac{\theta V_{1}^{\prime}(p_{1}(r))}{\Pi_{1}^{\prime\prime}(p_{1}(r))}-\frac{(1-\theta)V_{0}^{\prime}(p_{0}(r))}{\Pi_{0}^{\prime\prime}(p_{0}(r))}\right)\\ &=\left(\frac{-\Pi_{1}^{\prime\prime}(p_{1}(r)\Pi_{0}^{\prime\prime}(p_{0}(r)))}{\Pi_{0}^{\prime\prime}(p_{0}(r))+\Pi_{1}^{\prime\prime}(p_{1}(r))}\right)\left(\theta w_{1}(p_{1}(r))-(1-\theta)w_{0}(p_{0}(r))\right),\end{split}$ (17)
@@
Under Assumptions 4 and 5, $V_{\text{con}}(p^{*})>V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ if
@@
$\Pi^{\prime\prime}_{x}(p)=-2f_{x}(p)+f_{x}^{\prime}(p)(b-p).$
@@
Under Assumptions 4 and 5, $V_{\text{con}}(p^{*})<V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ if
@@
$\frac{\theta F_{1}(p^{*})/f_{1}(p^{*})}{2+L(p^{*})\alpha_{1}(p^{*})}<\frac{(1-\theta)F_{0}(p^{*})/f_{0}(p^{*})}{2+L(p^{*})\alpha_{0}(p^{*})}.$
F.3 Comparison of Decreasing Ratio Condition to Increasing Ratio Condition
The results in Section D.2.2 depend on the decreasing ratio condition (DRC) given in Assumption 5. This is analogous to the “increasing ratio condition (IRC)” from Aguirre et al. [2010], which says that the ratio $\frac{W_{x}^{\prime}(p)}{\Pi_{x}^{\prime\prime}(p)}$ is increasing in $p$. We now discuss in more detail the relationship between the DRC and the IRC, including sufficient conditions under which the DRC holds. In introducing the IRC, Aguirre et al. [2010] describe a “very large set of demand functions” for which the IRC holds. Aguirre et al. [2010] give sufficient conditions for the IRC to hold in Appendix B from their paper, which includes linear functions and exponential and constant elasticity functions.